OpenStudy (anonymous):

the 12th term of an arithmetic sequence is 36, and the common difference is 3. what is the first term of the sequence? a. 1 b. 2 c. 3 d. 4

OpenStudy (anonymous):

@Ethanxiaorox

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@AnimeAsgardian

OpenStudy (alexandervonhumboldt2):

3 i guess

OpenStudy (anonymous):

what about this one

OpenStudy (animeasgardian):

I agree, I think it's 3

OpenStudy (anonymous):

the common difference in an arithmetic sequence is -2 and the first term is 47. what is the 29th term? a. -9 b. -11 c. 11 d. 18

OpenStudy (alexandervonhumboldt2):

47, 45, 43, 41, .... let me solve

OpenStudy (anonymous):

thank you

OpenStudy (alexandervonhumboldt2):

47-(29*2)

OpenStudy (alexandervonhumboldt2):

-11

OpenStudy (anonymous):

what is the 13th term of the arithmetic sequence?\[\frac{ 3 }{ 4 },1,1\frac{ 1 }{ 4 },1\frac{ 1 }{ 2 }\] a. 3 b. 3 1/2 c. 4 d. 3 3/4

OpenStudy (alexandervonhumboldt2):

3/4+(1/4*13)

OpenStudy (alexandervonhumboldt2):

4

OpenStudy (anonymous):

i have to type this all back out cause i messed up -.-

OpenStudy (anonymous):

what is the general rule for the sequence? 1.11, 2.54, 3.97, 5.40 a. \[a_{n}=1.43n-0.32\] b. \[a_{n}=1.11n+1.43\] c. \[a_{n}=1.11n\] d. \[a_{n}=1+0.11n\]

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

B

OpenStudy (anonymous):

what is the 13th term of the following arithmetic sequence? 5,9,13,17 a. 52 b. 53 c. 54 d. 55

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@AnimeAsgardian

OpenStudy (alexandervonhumboldt2):

5+(13*4)

OpenStudy (alexandervonhumboldt2):

57

OpenStudy (anonymous):

57 isnt a choice ;o

OpenStudy (anonymous):

sorry i was away from my comp \(\sf a_n=a_1+(n-1)d\\a_{13}=5+(13-1)4\\a_{13}=?\)